OpenStudy (anonymous):

Here are 10 test scores: 50, 74, 76, 77, 78, 79, 80, 80, 82, 84. The mean of these scores is 76. How does removing the outlier 50 affect the mean? A. The set has 50 as an outlier and removing it decreases the mean by about 2. B. The set has 50 as an outlier and removing it increases the mean by about 3. C. The set has 50 as an outlier and removing it decreases the mean by about 6.

jimthompson5910 (jim_thompson5910):

take 50 out of the set and find the mean of 74, 76, 77, 78, 79, 80, 80, 82, 84 so add up those numbers and divide by 9 Mean = (74+76+77+78+79+80+80+82+84)/9 Mean = 710/9 Mean = 78.8888888888889 So we can see the mean goes from 76 to 78.8888888888889 when you take out the outlier of 50

OpenStudy (anonymous):

Calls the sum of the test scores S. S / 10 = 76 S = 760 Now, you take off the 50. (S - 50) / 9 = ? Change S to 760, you will have: (760-50)/9 = 710/9 = 78,89 It will increase about 3, like the letter B.

OpenStudy (anonymous):

thx